OpenStudy (anonymous):

HOMEWORK 10 anyone please!

OpenStudy (anonymous):

Please reply!!!!!!!!!!! ;)

OpenStudy (anonymous):

H10 P2 The impedance of circuit A, ZA=R1+R2+j*w*L+1/(j*w*C) As ω→0, ZA = inf As ω→∞, ZA = inf ZB= ??? As ω→0, ZB= 0 As ω→∞, ZB= 0 ZC=??? As ω→0, ZC= R2 As ω→∞, ZC= R1

OpenStudy (anonymous):

ZC =((R1+1/(j*w*C))*(R2+j*w*L))/((R1+1/(j*w*C))+(R2+j*w*L))

OpenStudy (anonymous):

ZB= (((j*w*L)*1/(j*w*C))/(j*w*L+(1/(j*w*C)))*(R1+R2))/(((j*w*L)*1/(j*w*C))/(j*w*L+(1/(j*w*C)))+(R1+R2)) that should just paste in... but if it doesn't "Parse" or gives some kind of error message... reference the attached pic.

OpenStudy (anonymous):

Thanx alot guys!!!!!!!!!!!! Do you know the correct equations H9P2's last two questions (1/R)*sqrt(L/C) 2*sqrt(L/C) these are wrong answers, plz let us know the correct one

OpenStudy (anonymous):

SergioAlexis NdeveR gratitude 4m my side.......... And plz don't forget to answer the question above.

OpenStudy (anonymous):

H9P2 Last two questions (R)*sqrt(C/L) 0.5*sqrt(L/C)

OpenStudy (anonymous):

H10P1 1. s*t 2. p*w 3. q*v 4. r*u

OpenStudy (anonymous):

@ olufemi H10P1 1. s*t 2. p*w 3. q*v 4. r*u H10P3 1) (((j*w*L+R2)*(1/(j*w*C))/(j*w*L+R2+(1/(j*w*C))))/((j*w*L + R2)*(1/(j*w*C))/(j*w*L+R2+(1/(j*w*C)))+R1)*(R2/(j*w*L + R2))) 2) ((j * w * L + R2) * (1 / (j * w * C))) / (j * w * L + R2 + (1 / (j * w * C ))) 3) (sqrt (R1/R2 -1)) / (w * R1) 4) ((sqrt (R1/R2 -1)) * R2) / w w = 2*pi*f 5) Cmatch = 1/(w*sqrt(R1R2 – R2^2)) - (formula) 6) Lmatch = sqrt(R1R2 – R2^2)/w - (formula)

OpenStudy (anonymous):

what do you mean by formula?

OpenStudy (anonymous):

for 5) and 6) part i hav written d formula coz everyone has diff. values so those who want can substitute their values and can reach to the final ans